(ii) Suppose X N(0,1) has the standard normal distribution. Find the probability distribution function ofY - 2
Answer #1

here as we know that pdf of X :f(x)=rac{1}{sqrt{2pi }}e^{-x^{2}/2}  

also as Y=1/X2

X =-/+ 1/Vy

|dx/dy| =(1/2)y-3/2

hence from transformation method:

pdf of Y f(y)=|dx/dy|*(fx(g(y))+fx(-g(y)))

=(1/2)y^{-3/2} *(rac{1}{sqrt{2pi }}e^{-(1/sqrt{y})^{2}/2}+rac{1}{sqrt{2pi }}e^{-(-1/sqrt{y})^{2}/2})

f(y)= y^{-3/2} *(rac{1}{sqrt{2pi }}e^{-rac{1}{2y}}) for 0 <y< infty